### v1

# Heun-function analysis of the Dirac spinor spectrum in a sine–Gordon soliton background

Heun-function analysis of the Dirac spinor spectrum in a sine–Gordon soliton background
H. Blas, R.P.N. Laeber Fleitas and J. Silva Barroso
a Instituto de Física
Universidade Federal de Mato Grosso
Av. Fernando Correa, $N^{0}$   2367
Bairro Boa Esperança, Cep 78060-900, Cuiabá - MT - Brazil.
We study the Dirac spectrum in a sine–Gordon soliton background, where the induced position-dependent mass reduces the spectral problem to a Heun-type differential equation. Bound and scattering sectors are treated within a unified framework, with spectral data encoded in Wronskians matching local Heun solutions and exhibiting explicit dependence on the soliton parameters and the bare fermion mass. This formulation enables a systematic analysis of spinor bound and scattering states, supported by analytic and numerical verification of wave function matching across the soliton domain. The present work is related to arXiv:2512.07658 and emphasizes a pedagogical treatment of scattering states within the Heun-equation formalism.
1 Introduction
Kink-fermion systems continue to attract significant interest due to their importance in nonperturbative quantum field theory [1]. These systems exhibit fermionic zero modes, charge fractionalization, and excited bound states, and are studied either via numerical approaches that include fermion back-reaction or within frameworks assuming a fixed kink background; for exact and self-consistent treatments of fermionic back-reaction, see [2, 3] and references therein.
On the other hand, Heun’s equation has become increasingly pivotal in modern theoretical and mathematical physics, offering a unifying framework for the analysis of a diverse range of complex phenomena that extend well beyond the capabilities of simpler special functions, such as the hypergeometric functions. It is the most general form of a second-order linear ordinary differential equation (ODE) possessing four regular singular points, thus generalizing the hypergeometric equation, which is limited to three singular points. This broad generality makes Heun’s equation a crucial tool in the study of physical systems that involve more intricate potentials, boundary conditions, or geometries, many of which naturally lead to equations of the Heun type [4, 5, 6, 7].
One of the key strengths of Heun’s equation lies in its ability to capture a wider variety of physical behaviors, including those associated with multi-scale interactions. In such cases, no single asymptotic expansion suffices to describe the system’s behavior across all scales; instead, a global analytic framework is required to connect disparate regions of the solution. Heun’s equation, with its multiple singularities, provides exactly such a global description, offering a unified approach to systems with complex, scale-dependent interactions.
Furthermore, Heun’s equation accommodates multi-pole interactions, which arise in physical scenarios where multiple distinct sources—each contributing a singularity—interact with one another. The presence of multiple singular points in the equation’s structure mirrors the contributions from these different physical sources, thus providing a natural mathematical framework for modeling such multi-pole interactions. As a result, Heun’s equation has found applications across a wide array of fields, from quantum mechanics and general relativity to condensed matter physics and string theory, where systems inherently exhibit multi-scale or multi-pole characteristics.
This connection between physical multiplicity (scales, poles, or sources) and analytic multiplicity (singular points) explains why Heun’s equation has become a universal descriptor of complex modern systems — from quantum fields around black holes to molecular and condensed-matter spectra. It captures the interference, coupling, and transition phenomena that occur when multiple characteristic structures coexist in a single system.
In quantum mechanics, Heun’s functions frequently appear in the study of: Anharmonic oscillators and double-well potentials. Two-center Coulomb problems (e.g., diatomic molecules). Quasi-exactly solvable systems, where a finite portion of the spectrum can be obtained analytically. The confluent forms of the Heun equation (e.g., confluent, biconfluent, double confluent) arise naturally when studying potentials such as the Mathieu, Razavy, or double sinh-Gordon potentials.
Heun equations are ubiquitous in modern gravitational and black hole physics. The Teukolsky equation, which governs perturbations of Kerr, Kerr–Newman, or Kerr–de Sitter spacetimes, can often be reduced to confluent or general Heun equations. The radial and angular parts of wave equations for scalar, electromagnetic, or gravitational fields in these geometries are Heun-type. This connection provides a bridge between exact analytical results and numerical relativity, relevant to quasi-normal modes and black hole spectroscopy.
In gauge/gravity duality, AdS/CFT correspondence and string theory, Heun-type equations arise in the analysis of field propagation in AdS black hole backgrounds, fluctuation spectra of D-branes, conformal blocks in 2D CFTs, where Heun functions encode the monodromy structure of conformal correlators. Thus, Heun functions serve as a link between differential geometry, conformal symmetry, and quantum field theory.
In condensed matter and optical physics, Heun equations describe graphene in external fields, photonic crystal modes, quasi-periodic potentials and Bloch waves, topological insulator edge states in nontrivial geometries. Their solutions describe localization, tunneling, and spectral band structures in systems with nonuniform or composite potentials.
The general Heun equation, the canonical local form with singularities at $0,1,a,\infty$, is written as
$H^{\prime\prime}(z)+(\frac{\gamma}{z}+\frac{\delta}{z-1}+\frac{\epsilon}{z-a})H^{\prime}(z)+\frac{\alpha\beta z-q}{z(z-1)(z-a)}H(z)=0,$ (1.1)
[…]
$\alpha+\beta+1=\gamma+\delta+\epsilon.$ (1.2)
[…]
From a mathematical physics perspective, Heun’s equation sits at the core of a hierarchy of Fuchsian equations. Its confluent limits define an entire taxonomy of special functions, each relevant to a distinct class of physical systems. Recent progress in symbolic computation (e.g., Maple, Mathematica) and numerical algorithms has made it feasible to use Heun functions in practical calculations, bridging the gap between exact analytic methods and computational physics.
In this paper we study the spectrum of a Dirac fermion in a sine–Gordon soliton background, where the induced position-dependent mass reduces the Dirac equation to a nontrivial eigenvalue problem with spinor components obeying Heun-type equations. This framework resolves the bound-state and scattering spectra in a unified manner, making explicit their dependence on the soliton parameters and bare fermion mass. We show that the Heun formalism is essential for constructing all nonzero-energy states, with the full scattering data encoded in Wronskians matching local Heun solutions across the soliton. The resulting spectral characterization provides a systematic basis for analyzing soliton–fermion stability and topologically protected excitations.
The paper is organized as follows. In the next section, we introduce the model describing a Dirac spinor coupled to a sine-Gordon soliton. Section 2.1 is devoted to the discussion of the Heun equation arising in the kink–spinor coupled system. In Section 3, we apply the Wronskian method to implement the matching of the scattering states at the origin. This section also analyzes the fermionic bound states and the phase shifts of the spinor components after propagation through the kink background. Finally, Section 4 presents the conclusions and discussions.
2 Dirac spinor coupled to a sine-Gordon soliton
When a Dirac spinor is coupled to a sine-Gordon soliton, the resulting fermion–soliton system exhibits a rich interplay between nonlinear and relativistic dynamics. The sine-Gordon soliton provides a localized, topologically nontrivial background field, while the Dirac spinor experiences this background as an effective potential with a nontrivial spatial structure. The coupled field equations will be reduced to a second-order differential equation governing the spinor components. These equations generally possess multiple singularities arising from the soliton profile and boundary conditions. Consequently, they are naturally expressible in terms of the Heun equation, the most general second-order linear ODE with four regular singular points. The emergence of the Heun structure reflects the multi-scale and multi-pole character of the fermion-soliton interaction, providing a unified analytic framework for exploring bound states, scattering, and spectral properties within this system.
Consider the following system of equations [3]
$(\partial_{t}+\partial_{x})\psi_{L}+Me^{2i\beta\varphi}\psi_{R}$ $=$ $0$ (2.1)
$(\partial_{t}-\partial_{x})\psi_{R}-Me^{-2i\beta\varphi}\,\psi_{L}$ $=$ $0,,$ (2.2)
where $\varphi$ is a sine-Gordon type soliton and $(\psi_{R},\psi_{L})^{T}$ is a two-component Dirac spinor. Let us define the parameterization
$\psi=e^{-iE_{1}t}\left(\begin{array}[]{c}u(x)\\ v(x)\end{array}\right),$ (2.5)
where the stationary spinor components $u(x)$ and $v(x)$ define the scattering solutions in the presence of the soliton $\varphi$, and $E_{1}$ is the energy of these states.
So, from (2.1)-(2.2) and (2.5) one can write the coupled system of static equations
$-E_{1}u+i\partial_{x}u+iMe^{-2i\beta\varphi}v$ $=$ $0,$ (2.6)
$E_{1}v+i\partial_{x}v+iMe^{2i\beta\varphi}u$ $=$ $0.$ (2.7)
We assume the next Ansatz for the scalar field as
$e^{2i\beta\varphi}=-(\frac{1+ie^{-2Kx}}{1-ie^{-2Kx}})^{2},$ (2.8)
with $K\in\hbox{\rm I\kern-1.79993ptR}$. The scalar function $\varphi(x)$ can be written as
$\varphi^{\pm}(x)=-\frac{2}{\beta}\arctan{\Big[e^{2Kx}\Big]},\,\,\,\,K\equiv\pm M.$ (2.9)
This is a kink(anti-kink). These functions exhibit the topological charges
$Q_{k(\bar{k})}$ $=$ $\frac{\beta}{2\pi}(\varphi^{\mp}(+\infty)-\varphi^{\mp}(-\infty)),\,\,\,\,\,k=kink,\,\,\,\bar{k}=antikink$ (2.10)
$=$ $\pm\frac{1}{2}.$ (2.11)
Therefore, the state with topological charge $Q_{kink-top}^{(I)}=+\frac{1}{2}$ is the kink and the state with $Q_{kink-top}^{(I)}=-\frac{1}{2}$ is the anti-kink. Notice that these charges are fractional, i.e. one-half of the integer $\pm 1$. In the Fig. 1 we plot this kink type soliton.
From (2.7) one can write
$v(x)=\frac{i}{M}(\frac{1+ie^{-2Kx}}{1-ie^{-2Kx}})^{2}(E_{1}u-iu^{\prime}(x)).$ (2.12)
Substituting this last relationship for $v$ into (2.6) one gets
$u^{\prime\prime}(x)-4iK\mbox{sech}(2Kx)u^{\prime}(x)+(E_{1}^{2}-M^{2}+4E_{1}K\mbox{sech}(2Kx))u(x)=0.$ (2.13)
Figure 1: (color online) Scalar kink (blue) for $M=1.5,\beta=1.$
2.1 Heun’s equation
In order to motivate the transformation of the equation (2.13) into a Heun-type equation, we can begin by analyzing the structure of this equation and identifying the key features that allow it to be recast in the form of the general Heun equation (1.1). To recast this equation as a Heun-type equation, we observe the following key features:
1. Singularities. The term $\mbox{sech}(2Kx)$ in (2.13) suggests the presence of singularities at $x=\pm\infty$; where the solution $u(x)$ will typically be localized around the soliton, and the equation will exhibit different asymptotic behavior at the boundaries of the soliton. Heun’s equation, in its general form, is known for its ability to model such situations with multiple regular singular points.
2. Non-constant coefficients. The presence of $\mbox{sech}(2Kx)$ as a coefficient in both the zero’th and first derivative terms introduces a form of “non-uniformity” into the equation, which is characteristic of Heun-type equations. In particular, Heun’s equation is capable of describing systems where the potential or interaction term varies spatially (or temporally), as is the case here with the soliton profile.
3. Change of variables. By introducing a new variable-typically a rescaling or reparametrization such as $y=-\tanh{(2Kx)}$ or another function that converts the $\mbox{sech}(2Kx)$ dependence into rational functions of $y$-the equation can be rewritten in a form with regular singularities located at finite points and, generally, an additional singularity at infinity. This transformation exposes the underlying Fuchsian character of the equation and places it naturally within the Heun family.
In this reformulated form, the equation inherits the characteristic analytic structure of Heun equations, including multiple singular points associated with the soliton-induced interaction scales. Recognizing the equation as Heun-type allows one to situate it within the broader analytic framework used to study systems with multi-scale potentials, nontrivial boundary behavior, and complex asymptotic structures. This connection enables the application of known Heun-function methods to analyze both local and global properties of the solutions.
4. Multi-scale structure and singularities. The localized soliton introduces distinct spatial scales, near its core and asymptotically far away, so no single asymptotic expansion can capture the solution across the entire domain. This multi-scale behavior, together with the presence of several singular points generated by the soliton profile and boundary conditions, naturally leads to a differential equation of Heun type. Mapping the problem to the general Heun equation accommodates the varying coefficients and singularity structure, allowing one to exploit established analytical tools for constructing and analyzing the fermion–soliton spectrum.
Firstly, we write this equation in a Fuchsian type second order ODE. So, as suggested above, let us make the next change of variable
$y=-\tanh{(2Kx)}$ (2.14)
in the differential equation for $u$ (2.13) in order to get
$u^{\prime\prime}(y)-2(\frac{y-i\sqrt{1-y^{2}}}{1-y^{2}})u^{\prime}(y)+\frac{E_{1}^{2}-M^{2}+4E_{1}K\sqrt{1-y^{2}}}{4K^{2}(1-y^{2})^{2}}u(y)=0.$ (2.15)
Next, remove the square root with the trigonometric substitution
$y=\cos{\theta}$ (2.16)
in order to get
$u^{\prime\prime}(\theta)+(\cot{\theta}-2i)u^{\prime}(\theta)+(\frac{E_{1}^{2}-M^{2}}{4K^{2}(\sin{\theta})^{2}}+\frac{E_{1}}{K\sin{\theta}})u(\theta)=0.$ (2.17)
Furthermore, rationalize with the substitution
$w=e^{i\theta}.$ (2.18)
So, one gets
$u^{\prime\prime}(w)+\frac{2}{w(w^{2}-1)}u^{\prime}(w)+\left(\frac{E_{1}^{2}-M^{2}}{K^{2}(w^{2}-1)^{2}}-(\frac{2iE_{1}}{K})\frac{1}{w(w^{2}-1)}\right)u(w)=0.$ (2.19)
This is a Fuchsian second order ODE with singular points at
$w=0,w=1,w=-1,w=\infty.$ (2.20)
Thus, one has four regular singular points, which is the defining feature of a Heun-type equation.
To match the standard Heun canonical locations one performs a linear (Möbius) map in the equation (2.19). We consider two cases.
The first case. Consider the next mapping
$z=\frac{w+1}{2}.$ (2.21)
This sends
$w=-1\rightarrow z=0,\,\,\,\,w=1\rightarrow z=1,\,\,\,\,w=0\rightarrow z=\frac{1}{2},\,\,\,\,w=\infty\rightarrow z=\infty.$ (2.22)
So, the third finite singularity is fixed at $a=\frac{1}{2}$. The equation (2.19) becomes
$u^{\prime\prime}(z)+[\frac{a}{z(z-1)(z-a)}]u^{\prime}(z)+[\frac{E_{1}^{2}-M^{2}}{4K^{2}z^{2}(z-1)^{2}}-(\frac{iE_{1}}{K})\frac{1}{z(z-1)(z-a)}]u(z)=0,\,\,\,\,a=\frac{1}{2}.$ (2.23)
Next, in order to write (2.23) in the canonical form of the Heun equation (1.1) one performs the usual gauge transformation as product of simple powers at the finite singularities. So, one has the transformation
$u(z)$ $=$ $f(z)\,H(z),$ (2.24)
$f(z)$ $=$ $z^{\mu}(z-1)^{\nu}(z-a)^{\sigma}.$ (2.25)
Let us write the identity
$u^{\prime\prime}(z)+Pu^{\prime}(z)+Qu(z)=f(z)[H^{\prime\prime}(z)+(P+2\frac{f^{\prime}}{f})H^{\prime}(z)+(Q+P\frac{f^{\prime}}{f}+\frac{f^{\prime\prime}}{f})H(z)],$ (2.26)
with
$P$ $=$ $\frac{a}{z(z-1)(z-a)}$ (2.27)
$=$ $\frac{1}{z}+\frac{1}{z-1}-\frac{1/a}{z-a}$
$Q$ $=$ $(\frac{E_{1}^{2}-M^{2}}{4K^{2}})\frac{1}{z^{2}(z-1)^{2}}-(\frac{iE_{1}}{K})[\frac{1}{z(z-1)(z-a)}],\,\,\,\,\,a=\frac{1}{2}.$ (2.28)
and
$\frac{f^{\prime}}{f}$ $=$ $\frac{\mu}{z}+\frac{\nu}{z-1}+\frac{\sigma}{z-a},$ (2.29)
$\frac{f^{\prime\prime}}{f}$ $=$ $(\frac{f^{\prime}}{f})^{\prime}+(\frac{f^{\prime}}{f})^{2}.$ (2.30)
Next, taking into account the relationship (2.27) and the coefficient of $u^{\prime}$ in (2.26), i.e. $(P+2\frac{f^{\prime}}{f})$, one can determine the parameters $\gamma,\delta,\epsilon$ appearing in the coefficient of $u^{\prime}$ of the standard Heun equation (1.1). So, one has
$\gamma=1+2\mu,\,\,\,\,\delta=1+2\nu,\,\,\,\,\,\epsilon=2\sigma-2.$ (2.31)
Now, let us consider the coefficient of $H(z)$ in (2.26). Taking into account (2.27)-(2.30) one can write
$Q+P\frac{f^{\prime}}{f}+\frac{f^{\prime\prime}}{f}$ $=$ $(\frac{E_{1}^{2}-M^{2}}{4K^{2}})\frac{1}{z^{2}(z-1)^{2}}-(\frac{iE_{1}}{K})[\frac{1}{z(z-1)(z-a)}]+$ (2.34)
$(\frac{1}{z}+\frac{1}{z-1}-\frac{2}{z-a})(\frac{\mu}{z}+\frac{\nu}{z-1}+\frac{\sigma}{z-a})+$
$(\frac{\mu}{z}+\frac{\nu}{z-1}+\frac{\sigma}{z-a})^{\prime}+(\frac{\mu}{z}+\frac{\nu}{z-1}+\frac{\sigma}{z-a})^{2}.$
Let us examine the double pole terms appearing in (2.34)-(2.34) at $z=0,1$ and $a$ and find the relevant parameters in order to convert them into simple pole expressions. So, after some simplifications, the terms with poles at $z=0$ and $1$ can be written as
$(\frac{E_{1}^{2}-M^{2}}{4K^{2}})\frac{1}{z^{2}(z-1)^{2}}+\frac{\mu^{2}}{z^{2}}+\frac{\nu^{2}}{(z-1)^{2}}$ $=$ $\frac{(\frac{E_{1}^{2}-M^{2}}{4K^{2}})+\mu^{2}-2\mu^{2}z+(\mu^{2}+\nu^{2})z^{2}}{z^{2}(z-1)^{2}}$ (2.35)
$=$ $-\frac{k^{2}}{2K^{2}}\frac{1}{z(z-1)},$ (2.36)
where the r.h.s in (2.36) with simple poles is achieved provided that
$E_{1}^{2}\equiv k^{2}+M^{2},\,\,\,\,\,\mu=-e_{1}\frac{ik}{2K},\,\,\,\,\nu=-e_{2}\frac{ik}{2K},\,\,\,\,\,e_{j}=\pm 1,\,j=1,2.$ (2.37)
In addition, the terms with double poles at $z=a$ become
$-\frac{3\sigma}{(z-a)^{2}}+\frac{\sigma^{2}}{(z-a)^{2}},$ (2.38)
whose total contribution vanishes provided that
$\sigma=0,3.$ (2.39)
Next, identify the $\alpha\beta$ and the accessory parameter $q$ in the coefficient of $H$. From (2.34)-(2.34) one can write the terms involving the simple poles as
$\frac{[-\frac{k^{2}}{2K^{2}}-(\mu+\nu)+2\sigma+2\mu\nu+2\mu\sigma+2\nu\sigma]\,z+\frac{3\mu}{2}-\frac{\nu}{2}-\mu\nu-\sigma-2\sigma\mu+\frac{k^{2}}{4K^{2}}-i\frac{E_{1}}{K}}{z(z-1)(z-a)}.$ (2.40)
So, comparing the numerator of (2.40) with the relevant term in (1.1), i.e. $(\alpha\beta z-q)$, one has
$\alpha\beta$ $=$ $-\frac{k^{2}}{2K^{2}}-(\mu+\nu)+2\mu\nu+2\sigma+2\mu\sigma+2\nu\sigma,$ (2.41)
$=$ $-(1+e_{1}e_{2})k\left(\frac{k-ie_{1}K}{2K^{2}}\right)+\left(-i\frac{(e_{1}+e_{2})k}{K}+2\right)\sigma$
$q$ $=$ $-\frac{k^{2}}{4K^{2}}+i\frac{E_{1}}{K}-\frac{3\mu}{2}+\frac{\nu}{2}+\mu\nu+\sigma+2\mu\sigma,$ (2.42)
$=$ $-\frac{(1+e_{1}e_{2})k^{2}}{4K^{2}}+ie_{1}\frac{(3-e_{1}e_{2})k}{4K}+\frac{iE_{1}}{K}+(1-ie_{1}\frac{k}{K})\sigma,$
$\alpha+\beta$ $=$ $2(\mu+\nu+\sigma)-1.$ (2.43)
$=$ $-\frac{ie_{1}(1+e_{1}e_{2})k}{K}+2\sigma-1.$
The last relationship (2.43) follows from (1.2) and (2.31). Then, one has
$\alpha$ $=$ $-\frac{1}{2}(1+e_{3}-2\sigma)-ie_{1}(1+e_{1}e_{2})\frac{k}{2K};\,\,\,\,\,\,\,\,\,e_{a}=\pm 1;\,\,\,a=1,2,3.$ (2.44)
$\beta$ $=$ $-\frac{1}{2}(1-e_{3}-2\sigma)-ie_{1}(1+e_{1}e_{2})\frac{k}{2K}.$ (2.45)
For the both values of the product $e_{1}e_{2}=\pm 1$ in (2.37), respectively, one can write the next sets of parameters
$e_{1}e_{2}=1,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\alpha$ $=$ $-\frac{1}{2}(1+e_{3}-2\sigma)-e_{1}\frac{ik}{K},\,\,\,\,\,\,\,\beta=-\frac{1}{2}(1-e_{3}-2\sigma)-e_{1}\frac{ik}{K},$ (2.46)
$\nu$ $=$ $\mu=-e_{1}\frac{ik}{2K},\,\,\,\,\gamma=\delta=1-e_{1}\frac{ik}{K},$ (2.47)
$q$ $=$ $-\frac{k^{2}}{2K^{2}}+i\frac{2E_{1}+e_{1}k}{2K}+(1-ie_{1}\frac{k}{K})\sigma;$ (2.48)
$e_{1}e_{2}=-1,\,\,\,\,\,\,\,\,\,\alpha$ $=$ $-\frac{1}{2}(1+e_{3}-2\sigma),\,\,\,\,\,\,\,\beta=-\frac{1}{2}(1-e_{3}-2\sigma)\,\,\,\,\,\,\,\,\,\,\,\mu=-\nu=-e_{1}\frac{ik}{2K},$ (2.49)
$\gamma$ $=$ $1-e_{1}\frac{ik}{K},\,\,\,\,\,\,\,\,\delta=1+e_{1}\frac{ik}{K},$ (2.50)
$q$ $=$ $i\frac{E_{1}+e_{1}k}{K};$ (2.51)
In the following we consider the second case $e_{1}e_{2}=-1$ above, i.e. the Eqs. (2.49)-(2.51). Let us consider the next parameter values
$\sigma=0,\,\,\,e_{3}=1,\,\,\,\,e_{1}=\pm 1.$ (2.52)
Then one has
$\alpha$ $=$ $-1,\,\,\,\,\beta=0,\,\,\,\,\,q=i\frac{E_{1}+e_{1}k}{K}$ (2.53)
$\gamma$ $=$ $\delta^{\star}=1-e_{1}\frac{ik}{K},\,\,\,\,\mu=-\nu=-e_{1}\frac{ik}{2K}$ (2.54)
Therefore, the spinor components $u(z)$, which follows from (2.24)-(2.25), can be written as
$u_{1}(z)$ $=$ $z^{\mu}(z-1)^{\nu}Hl[a,q;\alpha,\beta,\gamma,\delta;z]$ (2.55)
$z$ $=$ $\frac{1}{1-ie^{2Kx}},$ (2.56)
where $Hl[a,q;\alpha,\beta,\gamma,\delta;z]$ denotes the six-parameter local Heun function [4, 7, 5]. The variable $z$ in (2.56) follows after the successive variable transformations (2.14), (2.16), (2.18), and (2.21) are performed, i.e. $x\rightarrow y$, $y\rightarrow\theta$, $\theta\rightarrow w$ and $w\rightarrow z$, respectively.
In addition, performing similar successive variable transformations, the $v_{1}(z)$ component can be written from the equation (2.12) as
$v_{1}(z)$ $=$ $-\frac{i}{4M}\frac{1}{(z-a)^{2}}\left(iE_{1}u_{1}(z)+2Kz(z-1)u_{1}^{\prime}(z)\right),$ (2.57)
where $u_{1}(z)$ is provided in (2.55).
For a fermionic mode incident on the kink from the left, let us consider the solution of Eq. (2.13) in terms of the Heun’s local function $Hl[a,q;\alpha,\beta,\gamma,\delta;z]$, such that asymptotically it must reduce to the transmitted plane wave, being $\sim e^{ikx}$. Then, the corresponding solution becomes
$u_{1}^{(1)}(x)$ $=$ $z^{\mu}(z-1)^{\nu}Hl[a,q;\alpha,\beta,\gamma,\delta;z],\,\,\,\mu=-\nu=\frac{ik}{2K},$ (2.58)
$=$ $e^{\frac{\pi k}{4K}}\,e^{ikx}\,Hl[\frac{1}{2},i\frac{E_{1}+k}{K};-1,0,1-i\frac{k}{K},1+i\frac{k}{K};\frac{1}{1-ie^{2Kx}}],$ (2.59)
where the notation $Hl[\frac{1}{2},q;\alpha,\beta,\gamma,\delta;z]$ is used for the six parameter local Heun function [4, 5]. Note that we have used the parameter set provided in (2.52)-(2.54) for $e_{1}e_{2}=-1,e_{1}=1.$ In fact, the behavior of $z$ in the original spatial coordinate becomes $z\rightarrow ie^{-2Kx}$ in the limit $x\rightarrow+\infty$. So, the argument of the local Heun function tend to zero (one of the regular singular points at $z=0$) as $x\rightarrow+\infty$. The local Heun function $Hl[\frac{1}{2},q;\alpha,\beta,\gamma,\delta;z]$ is analytic at the regular singular point $z=0$ and is normalized such that $Hl(0)=1$. Around this point, it admits a convergent Taylor expansion. In the complex $z$-plane the radius of convergence of this local series is given by $min\{|a|,1\}$; for $a=1/2$, this yields a convergence radius of $1/2$.
The second case. Consider the next transformation in the Fuchsian equation (2.19)
$z=\frac{1-w}{2}.$ (2.60)
This sends
$w=-1\rightarrow z=1,\,\,\,\,w=1\rightarrow z=0,\,\,\,\,w=0\rightarrow z=1/2,\,\,\,\,w=-\infty\rightarrow z=\infty.$ (2.61)
So, the third finite singularity is also fixed at $a=\frac{1}{2}$. The equation (2.19) becomes
$u_{2}^{\prime\prime}(z)+[\frac{a}{z(z-1)(z-a)}]u_{2}^{\prime}(z)+[\frac{E_{1}^{2}-M^{2}}{4K^{2}z^{2}(z-1)^{2}}+(\frac{iE_{1}}{K})\frac{1}{z(z-1)(z-a)}]u_{2}(z)=0,\,\,\,\,a=\frac{1}{2}.$ (2.62)
Remarkably, this equation is similar to the first case equation in (2.23) except that the sign of the term with the factor $(\frac{iE_{1}}{K})$ is reversed. Therefore, in order to write (2.62) in the canonical form (1.1) one performs similar gauge transformation as product of simple powers at the finite singularities as in (2.24)-(2.25). Following similar for the case $e_{1}e_{2}=-1$ above, we can get the next set of parameters
$\sigma$ $=$ $0,\,\,\,e_{3}=1,\,\,\,\,e_{1}=\pm 1.$ (2.63)
$\alpha$ $=$ $-1,\,\,\,\,\beta=0,\,\,\,\,\,q=-i\frac{E_{1}+e_{1}k}{K}$ (2.64)
$\gamma$ $=$ $\delta^{\star}=1+e_{1}\frac{ik}{K},\,\,\,\,\mu=-\nu=e_{1}\frac{ik}{2K}.$ (2.65)
So, in this case the spinor component $u(z)$ being a solution of (2.62), which follows from (2.24)-(2.25), can be written as
$u_{2}(z)$ $=$ $z^{\mu}(z-1)^{\nu}Hl[\frac{1}{2},-i\frac{E_{1}+k}{K};-1,0,1+\frac{ik}{K},1-\frac{ik}{K};z]$ (2.66)
$z$ $=$ $\frac{1}{1+ie^{-2Kx}},$ (2.67)
where the parameters were taken from (2.63)-(2.65) with $e_{1}=1$.
In addition, performing similar successive variable transformations, the $v(z)$ component can be written from the equation (2.12) as
$v_{2}(z)$ $=$ $-\frac{i}{4M}\frac{1}{(z-a)^{2}}\left(iE_{1}u_{2}(z)-2Kz(z-1)u_{2}^{\prime}(z)\right),$ (2.68)
where $u_{2}(z)$ is provided in (2.66)-(2.67).
Some comments are in order here. First, the behavior of the both types of solutions above are different at the asymptotic values $x\rightarrow\pm\infty$ corresponding to the regular singularities $z=0,1$. The argument of $u_{1}(z)$ tends to zero as $x\rightarrow\infty$, whereas the argument of $u_{2}(z)$ tends to zero as $x\rightarrow$-$\infty$.
Second, in order to get the transmission and reflection coefficients, it is needed the asymptotic behavior of $u_{1}$ as $x\rightarrow$-$\infty$. However, the Heun function $Hl[a,q;\alpha,\beta,\gamma,\delta;z]$ is only analytical in the range of $|z|<min\{a,1\}$. As $x\rightarrow$-$\infty$, we have that the argument of $u_{1}$ behaves as $z=\frac{1}{1-ie^{2Kx}}\rightarrow 1$. So, $u_{1}$ is thus not analytical [4, 7, 5]. This problem can be solved by considering another solution with good asymptotic behavior as $x\rightarrow$-$\infty$ [9, 10, 11]. The second type solution $u_{2}$ meets this condition, since its argument tends to zero as $x\rightarrow$-$\infty$, i.e. $z=\frac{1}{1+ie^{-2Kx}}\rightarrow 0$. As we will see below, using the both type of solutions $u_{1}$ and $u_{2}$ obtained in this way and the matching conditions for the fermionic wave functions, one can derive general expressions for the transmission and reflection coefficients.
Third, if $\gamma\neq\hbox{\sf Z\kern-3.99994ptZ}$, the Heun equation has two linearly independent solutions near $z=0$ [4, 7, 5]. From the relationship (2.24)-(2.25), since $\frac{k}{K}$ is real for the scattering states, one can get two linearly independent solutions $u_{1}^{(1)}(z)$ and $u_{1}^{(2)}(z)$ as
$u_{1}^{(1)}(z)$ $=$ $z^{\mu}(z-1)^{\nu}Hl[a,q;\alpha,\beta,\gamma,\delta;z]$ (2.69)
$u_{1}^{(2)}(z)$ $=$ $z^{\mu}(z-1)^{\nu}z^{1-\gamma}Hl[a,q+(\epsilon+\delta a)(1-\gamma);\alpha-\gamma+1,\beta-\gamma+1,2-\gamma,\delta;z],$ (2.70)
where the parameters (2.52)-(2.54) can be used.
A similar construction can be performed for the second case, providing two linearly independent solutions $u_{2}^{(1)}(z)$ and $u_{2}^{(2)}(z)$ with parameters (2.63)-(2.65). In fact, for the second type of solutions the argument of the local Heun function tend to zero as $x\rightarrow$-$\infty$, meaning that the local Heun function remains analytic in this limit. Note that the behavior of $z$ in the original spatial coordinate becomes $z\rightarrow-ie^{2Kx}$ from (2.67). Then, a first solution of (2.62) from (2.66)-(2.67) can be written as
$u_{2}^{(1)}(x)$ $=$ $z^{\mu}(z-1)^{\nu}\,Hl[\frac{1}{2},-i\frac{E+k}{K};-1,0,1+i\frac{k}{K},1-i\frac{k}{K};\frac{1}{1+ie^{-2Kx}}],$ (2.71)
$=$ $e^{-\frac{\pi k}{4K}}e^{ikx}Hl[\frac{1}{2},-i\frac{E+k}{K};-1,0,1+i\frac{k}{K},1-i\frac{k}{K};\frac{1}{1+ie^{-2Kx}}]$ (2.72)
where the parameters correspond to the set (2.63)-(2.65) with $e_{1}=1$.
The second solution becomes
$u_{2}^{(2)}(x)$ $=$ $z^{\mu}(z-1)^{\nu}z^{1-\gamma}Hl[a,q+(a\delta+\epsilon)(1-\gamma);\alpha+1-\gamma,\beta+1-\gamma,2-\gamma,\delta;\frac{1}{1+ie^{-2Kx}}],$ (2.74)
$=$ $e^{-\frac{\pi k}{2K}}e^{-ikx}(i+e^{2Kx})^{\frac{ik}{2K}}\times$
$Hl[\frac{1}{2},-i\frac{2E_{1}-k}{2K}-\frac{k^{2}}{2K^{2}};-1-\frac{ik}{K},-\frac{ik}{K},1-\frac{ik}{K},1-\frac{ik}{K};\frac{1}{1+ie^{-2Kx}}],$
Fourth, in (2.69)-(2.70) the Heun function is an infinite series, $Hl[a,q;\alpha,\beta,\gamma,\delta;z]=\sum_{n=0}^{\infty}h_{n}z^{n}$. The coefficients $h_{n}$ are determined by the three-term recurrence relationship
$R_{n-1}h_{n-1}+P_{n}h_{n}+Q_{n+1}h_{n+1}=0,$ (2.75)
such that the initial conditions $h_{0}=1$ and $h_{-1}=0$ are assumed. Here $R_{n}=(n+\alpha)(n+\beta),P_{n}=-q-n(n-1+\gamma)(1+a)-n(a\delta+\epsilon)$, and $Q_{n}=an(n-1+\gamma)$.
Fifth, in addition to solutions expressed in terms of Heun functions, exact closed-form solutions in elementary functions also exist under specific conditions, as examined below. It has been demonstrated that, for certain special parameter regimes
$\alpha,\beta$ $=$ $-N,\,\,\,\,\,N=0,1,2,$... (2.76)
$h_{N+1}$ $=$ $0,$ (2.77)
the Heun function $Hl[a,q;\alpha,\beta,\gamma,\delta;z]$ admits a finite-series representation and truncates to a polynomial in $z$ [4, 5]. This truncated condition allow us to obtain exact analytical solutions. In fact, in our case above for the solutions $u_{1}^{(1)}(z)$ and $u_{2}^{(1)}(z)$ one has $\alpha=-1$; so, the condition $h_{2}=0$ terminates those series up to the first order in $z$.
3 Wronskian method and matching scattering states
The solutions to the differential equation (2.13) can be expressed in terms of the local Heun functions [4, 7, 5]. Note that we have used the parameter set provided in (2.52)-(2.54) for $e_{1}e_{2}=-1,e_{1}=1.$ In Eq. (2.59), the argument of the local Heun function tend to zero (one of the regular singular points at $z=0$) as $x\rightarrow+\infty$. The local Heun function $Hl[\frac{1}{2},q;\alpha,\beta,\gamma,\delta;z]$ is analytic at the regular singular point $z=0$ and is normalized such that $Hl(0)=1$. Around this point, it admits a convergent Taylor expansion. In the complex $z$-plane the radius of convergence of this local series is given by $min\{|a|,1\}$; for $a=1/2$, this yields a convergence radius of $1/2$.
Although the Taylor expansion of $Hl[\frac{1}{2},q;\alpha,\beta,\gamma,\delta;z]$ possesses only a finite radius of convergence, the function itself admits analytic continuation to the entire complex plane, with a branch cut conventionally taken along $\frac{1}{2},\infty]$. It is thus well-defined at all finite points of the complex plane except at the regular singularities $z=1/2$ and $z=1$. In the Eq. (2.59), the argument of the local Heun function approaches unity as $x\rightarrow$-$\infty$, implying that these local solutions cease to provide a valid representation in this asymptotic regime.
However, as discussed above, for the second type of solutions the argument of the local Heun function tend to zero as $x\rightarrow$-$\infty$. Therefore, the general solution will be a linear combination of (2.72) and (2.74)
$u_{s}(x)=c_{1}u_{2}^{(1)}(x)+c_{2}u_{2}^{(2)}(x).$ (3.1)
The solutions given in Eqs. (2.59), $u_{1}^{(1)}$, and $u_{s}$ in (3.1) possess overlapping domains of analyticity in the variable $x$ . By equating these representations $u_{1}^{(1)}|_{x=x_{0}}=u_{s}|_{x=x_{0}}$ together with their first derivatives $u_{1}^{(1)^{\prime}}|_{x=x_{0}}=u^{\prime}_{s}|_{x=x_{0}}$, one can determine the coefficients $c_{1}$ and $c_{2}$ appearing in Eq. (3.1). The specific choice of the matching point $x=x_{0}$ is immaterial, provided it lies within the interval $(-\infty,+\infty)$. For convenience and by symmetry, we select $x_{0}=0$ as the matching point. So, one has
$u_{1}^{(1)}|_{(x=0)}$ $=$ $c_{1}u_{2}^{(1)}|_{(x=0)}+c_{2}u_{2}^{(2)}|_{(x=0)},$ (3.2)
$\frac{d}{dx}u_{1}^{(1)}|_{(x=0)}$ $=$ $c_{1}\frac{d}{dx}u_{2}^{(1)}|_{(x=0)}+c_{2}\frac{d}{dx}u_{2}^{(2)}|_{(x=0)}.$ (3.3)
This system provides an explicit expressions
$c_{1}=\frac{W(u_{1}^{(1)},u_{2}^{(2)})}{W(u_{2}^{(1)},u_{2}^{(2)})}|_{(x=0)},\,\,\,\,\,\,\,\,c_{2}=-\frac{W(u_{1}^{(1)},u_{2}^{(1)})}{W(u_{2}^{(1)},u_{2}^{(2)})}|_{(x=0)},$ (3.4)
where $W(v,w)\equiv vw^{\prime}-wv^{\prime}$ is the Wronskian of the two functions $v$ and $w$.
Figures 2, 3, 4, 5 and 6 display the real and imaginary parts of the scattering states corresponding to the spinor components, for $\beta=1,M=5,k=2.5,E_{1}=+5.6$. These plots qualitatively demonstrate the effectiveness of the matching conditions imposed at $x=0$ for the relevant scattering solutions. In particular, they illustrate the role of the Wronskian method in determining the coefficients $c_{1}$ and $c_{2}$ appearing in Eq. (3.4).
The Figs. 2 show the individual components around $x=0$ of the $u$-component. Scalar kink (blue) in the left and right Figs. The real and imaginary components of the incident, reflected and transmitted scattering wave function $u(x)$: Left Fig. $Re[c_{1}u_{in}(x)]$ (green), $Re[c_{2}u_{ref}(x)]$ (brown), $Re[u_{tr}(x)]$ (red). The right Fig. shows $Im[c_{1}u_{in}(x)]$ (green) and $Im[c_{2}u_{ref}(x)]$ (brown), $Im[u_{tr}(x)]$ (red).
In order to inspect their behavior around $x=0$ we plot the Figs. 3, 4, 5, and 6. Figure 3 depicts the scalar kink (blue) localized around the origin. The real part of the scattering wave function $u(x)$ in the region $x<0$ is shown as the combination
$Re[c_{1}u_{in}(x)+c_{2}u_{ref}(x)],$ (3.5)
with green line, while in the region $x>0$ it is plotted as
$Re[u_{tr}(x)],$ (3.6)
with red line. The inset highlights the smooth and continuous behavior of the $u$-component at the origin. This provides qualitative confirmation of the effectiveness of the Wronskian-based matching procedure imposed at $x=0$.
Furthermore, figure 4 depicts the same scalar kink (blue) localized around the origin. The imaginary part of the scattering wave function $u(x)$ in the region $x<0$ is shown as the combination
$Im[c_{1}u_{in}(x)+c_{2}u_{ref}(x)],$ (3.7)
with magenta line, while in the region $x>0$ it is plotted as
$Re[u_{tr}(x)],$ (3.8)
with red line. The inset emphasizes the smooth and continuous behavior of the imaginary component of $u$ at the origin. This qualitatively confirms the effectiveness of the Wronskian-based matching procedure applied at $x=0$.
The Figs. 5 and 6 show the real and imaginary parts of the $v$-component. These figures show qualitatively similar behavior as the $u$-component described above.
3.1 Fermionic bound states
We now turn to the fermionic bound states associated with the kink background, employing both the Heun-function and Heun-polynomial approaches. An analysis of the asymptotic behavior of Eq. (2.13) in the limit $|x|\rightarrow+\infty$ shows that the continuum (scattering) spectrum is characterized by
$E_{1}^{2}=M^{2}+k^{2},\,\,\,\,k\in\hbox{\rm I\kern-1.79993ptR},$ (3.9)
and the bound-state spectrum satisfies
$E_{bs}^{2}=M^{2}-\kappa^{2},\,\,\,\,\kappa>0.$ (3.10)
From a spectral perspective, bound states correspond to a purely imaginary continuation of the quasi-momentum into the complex plane, $k\rightarrow i\kappa$, and therefore appear as isolated eigenvalues lying above and below the continuum threshold at $\pm M$. Accordingly, Eqs. (2.12) and (2.13) imply that the $u$ and $v$ components of a fermionic bound state decay as $e^{-\kappa|x|}$ as $|x|\rightarrow+\infty$.
To analyze the bound states, we examine the behavior of the scattering solutions under the analytic continuation $k\rightarrow i\kappa$. In Eq. (2.58), the argument of the local Heun function tends to zero as $x\rightarrow+\infty$, causing the function to approach unity. Consequently, after the substitution $k\rightarrow i\kappa$ the transmitted fermionic wave exhibits the correct bound-state asymptotic behavior, decaying as $e^{-\kappa x}$ as $x\rightarrow+\infty$.
Similarly, replacing $k$ by $i\kappa$ in Eq. (2.74) yields a reflected wave with the proper bound-state asymptotics $e^{\kappa x}$ as $x\rightarrow$-$\infty$. However, the incident wave transforms into a term proportional to $e^{-\kappa x}$, which diverges in this limit. To eliminate this unphysical behavior of the incident fermionic component, the coefficient $c_{1}(E_{1},k)$ must vanish at $E_{1}=E_{1n}$, $k=i\kappa=i\sqrt{M^{2}-E_{1n}}$, where $E_{1n}$ denotes the bound state energy.
Since the coefficient $c_{1}(E_{1},k)$ is explicitly given by Eq. (3.4), the fermionic bound-state energy levels $E_{1n}$ are determined by the solutions of the resulting transcendental equation
$c_{1}(E_{1n},i\sqrt{M^{2}-E_{1n}})=0,\,\,\,n=0,1,2,$... (3.11)
3.2 Phase shifts and Levinson’s theorem
The coefficients $c_{1}$ and $c_{2}$ encapsulate the complete information regarding the scattering of the fermionic wave by the sine-Gordon type kink. From Eqs. (2.59) and (3.1), it follows that the asymptotic behavior of the $u$-component of the spinor wavefunction can be schematically expressed as
$c_{1}e^{ikx}\rightarrow e^{ikx}+c_{2}e^{\frac{\pi k}{4K}}e^{-ikx},$ (3.12)
which describes the decomposition of the incident wave into its transmitted and reflected components.
During the scattering process, the transmitted fermionic wave component $u$ acquires a phase shift $\delta_{u}$ relative to the incident wave. According to Eq. (3.12), this phase shift is given by
$\delta_{u}(k)=-\arg{[c_{1}(k)]}.$ (3.13)
Figure 2: (color online) Scalar kink (blue) in the left and right Figs. The real and imaginary parts of the incident, reflected and transmitted scattering wave function $u(x)$: Left Fig. $Re[c_{1}u_{in}(x)]$ (green), $Re[c_{2}u_{ref}(x)]$ (brown), $Re[u_{tr}(x)]$ (red), and right Fig. $Im[c_{1}u_{in}(x)]$ (green) and $Im[c_{2}u_{ref}(x)]$ (brown), $Im[u_{tr}(x)]$ (red). For $\beta=1,M=5,k=2.5,E_{1}=+5.6$.
Figure 3: (color online) Scalar kink (blue) around $x=0$. The real part of the scattering wave function $u(x)$ for $x<0$, $Re[c_{1}u_{in}(x)+c_{2}u_{ref}(x)]$ (green), and for $x>0$, $Re[u_{tr}(x)]$ (red). The inset figure shows the smooth matching at $x=0$.
Figure 4: (color online) Scalar kink (blue) around $x=0$. The imaginary part of the scattering wave function $u(x)$ for $x<0$, $Re[c_{1}u_{in}(x)+c_{2}u_{ref}(x)]$ (green), and for $x>0$, $Im[u_{tr}(x)]$ (red). The inset figure shows the smooth matching at $x=0$.
Figure 5: (color online) Scalar kink (blue) around $x=0$. The real part of the scattering wave function $v(x)$ for $x<0$, $Re[c_{1}v_{in}(x)+c_{2}v_{ref}(x)]$ (green), and for $x>0$, $Re[v_{tr}(x)]$ (red). The inset figure shows the smooth matching at $x=0$.
Figure 6: (color online) Scalar kink (blue) around $x=0$. The imaginary part of the scattering wave function $v(x)$ for $x<0$, $Re[c_{1}v_{in}(x)+c_{2}v_{ref}(x)]$ (green), and for $x>0$, $Re[v_{tr}(x)]$ (red). The inset figure shows the smooth matching at $x=0$.
Figure 7: (color online) Phase shift $\delta(k)$ vs $k$ of the scattering states $u$ and $v$ for the fermion mass $M=2.15\times 10^{-5}$. Note that $\delta(0)-\delta(+\infty)=\frac{\pi}{2}$.
In a similar manner, combining Eqs. (2.59) and (3.1) and accounting for Eq. (2.12), one finds that the asymptotic behavior of the $v$-component of the spinor wave function can be schematically expressed as
$c_{1}e^{ikx}\rightarrow e^{ikx}+c_{2}(\frac{E_{1}-k}{E_{1}+k})\,e^{\frac{\pi k}{2M}}e^{-ikx},$ (3.14)
such that the incident wave is decomposed into its transmitted and reflected components.
Consequently, the transmitted fermionic $v$-component acquires a phase shift $\delta_{v}(k)$ with respect to the incident wave. From Eq. (3.14), this phase shift is given by
$\delta_{v}(k)=-\arg{[c_{1}(k)]}.$ (3.15)
One therefore concludes that the upper and lower components of the spinor acquire identical phase shifts after scattering off the sine-Gordon–type soliton. This result contrasts with the recent findings in the ATM model with a soliton background of variable topological charge, where the phase shifts of the $u$ and $v$ components differ by a constant. In that case, a single phase shift is defined as the average of the phase shifts of the upper and lower components [2]. In the Fig. 5 we plot the common phase shift $\delta(k)=\delta_{u}(k)=\delta_{v}(k)$. Using this Fig. one can check the Levinson’s theorem [12]
$\delta(0)-\delta(+\infty)=\pi(n_{b}-\frac{1}{2}),$ (3.16)
where $n_{b}$ is the number of bound states in a given scattering channel. One notices that the Levinson’s theorem is satisfied for $n_{b}=1$. In the recent contribution [3] it has been demonstrated that the transcendental equation (3.11) supports the bound state energies $E_{10}=0$ and $E_{10}=\pm 0.8M$, which are the zero and valence modes. This shows that in the positive energy channel +0.8M+0.8M there is a bound state $n_{b}=1$.
4 Discussions
In this work, we have investigated the scattering and bound-state properties of a Dirac spinor coupled to a sine-Gordon soliton background. By formulating the spinor equations in the presence of the kink profile, we showed that the resulting differential equations for the spinor components can be expressed in terms of Heun-type equations (1.1). This formulation provides an exact and systematic framework for analyzing the fermionic spectrum in a nontrivial topological background.
The scattering problem was addressed through the implementation of the Wronskian method, which allowed us to impose consistent matching conditions at the origin (3.2)-(3.3). This approach proved to be particularly effective in determining the scattering coefficients (3.4) and ensuring the smooth behavior of the spinor components across the kink. The numerical results for the real and imaginary parts of the scattering states qualitatively confirm the validity and robustness of the Wronskian-based matching procedure (see the Figs. 2, 3, … 6).
Furthermore, we analyzed the fermionic bound states by means of analytic continuation of the scattering solutions, $k\rightarrow i\kappa$. The bound-state energies were shown to be determined by the zeros of a transcendental equation arising from the vanishing of the incident-wave coefficient (3.11). This condition guarantees the correct asymptotic behavior of the bound-state wave functions and leads to a discrete spectrum of fermionic energy levels localized around the kink.
In addition, we examined the phase shifts acquired by the spinor components during scattering off the kink background. These phase shifts encode important information about the interaction between the fermionic field and the topological soliton and may be relevant for understanding transport and polarization effects in related field-theoretical and condensed-matter systems, see [3] and references therein.
Heun’s equation serves as a universal mathematical structure in modern theoretical physics, encompassing phenomena from quantum to gravitational scales. It captures the essential complexity of real physical systems that lie beyond the reach of classical hypergeometric analysis and provides a powerful analytic framework for exploring nontrivial geometries, potentials, and symmetries.
The methods and results presented here can be extended in several directions. Possible future investigations include the study of other solitonic backgrounds, the incorporation of additional interaction terms, or the analysis of finite-temperature and time-dependent effects. The present framework may also find applications in effective models of low-dimensional systems where fermions interact with topological defects, as in the Ref. [13].